Hookworms, parasitic nematodes transmitted through contact between bare feet and soil, infect nearly a half billion people. In the small SE Asian nation of Timor-Leste, infection rates in some regions approach 70%, and the resulting anemia from these infections has a serious impact on quality of life. Health officials are concerned by a recent rise in Ivermectin-resistant cases in the isolated enclave of Oecussi, and are worried about these drug resistance genes spreading to the capital Dili, 165 km away. You have been tasked with conducting a population genetic survey to assess this risk. You genotype diploid female hookworms, Ancylostoma ceylanicum, from 100 patients in each location for a neutral locus with two alleles A1 & A2.

You find the following genotypic count data:

A1/A1

A1/A2

A2/A2

Oecussi

36

48

16

Dili

9

42

49

4a.Use these data to calculate FSTbetween these two populations.

4b. Assuming that these two populations are in drift/migration equilibrium, what is the effective number of migrants [*note –we ‘ll get to this on Thursday]moving between them each generation.
Answer #1

4a. The is a formula to find Fst.

F st = HT - HS - HT

Where HT is the heterozygosity in both population and HS is the heterozygosity in one population

H is calculated easely, it is 1 minus the summatory of the squared allele frequencies.

Η = 1-Σ2

Let's calculate allele frequencies:

In Oecussi.

p = 2(36)+48 / 200 = 0.6

q = 1 - p = 1 - 0.6 = 0.4

In Dili.

p = 2(9)+42 / 200 = 0.3

q = 1 - p = 1 - 0.3 = 0.7

For both populations together allele frequencies are:

p = 2(45)+90 / 400 = 0.45

q = 1 - p = 1 - 0.3 = 0.55

Now let's calculate HT.

HT = 1 - ((0.45)2+(0.55)2) = 1 - 0.505 = 0.495

Now for HS.

HS = 1 - ((0.6)2+(0.4)2) = 1 - 0.52 = 0.48

Now for Fst

Fst = HT - HS / HT = 0.495 - 0.48 / 0.495 = 0.03. A low value here indicates populations have a high similarity.

4b. There is formula that relates Fst to Nem (effective migrants).

Fst = 1+ 4Nem

If we clear the formula for Nem we have:

Nem = (1/Fst) - 1 -= 8.08

8.08 effective migrants